# Preference Elicitation for Offline Reinforcement Learning

### v1

@@ Abstract
Applying reinforcement learning (RL) to real-world problems is often made challenging by the inability to interact with the environment and the difficulty of designing reward functions. Offline RL addresses the first challenge by considering access to an offline dataset of environment interactions labeled by the reward function. In contrast, Preference-based RL does not assume access to the reward function and learns it from preferences, but typically requires an online interaction with the environment. We bridge the gap between these frameworks by exploring efficient methods for acquiring preference feedback in a fully offline setup. We propose Sim-OPRL, an offline preference-based reinforcement learning algorithm, which leverages a learned environment model to elicit preference feedback on simulated rollouts. Drawing on insights from both the offline RL and the preference-based RL literature, our algorithm employs a pessimistic approach for out-of-distribution data, and an optimistic approach for acquiring informative preferences about the optimal policy. We provide theoretical guarantees regarding the sample complexity of our approach, dependent on how well the offline data covers the optimal policy. Finally, we demonstrate the empirical performance of Sim-OPRL in different environments.
@@ 1 Introduction
Our contributions are the following: (1) In Section 3, we first formalize the new problem setting of preference elicitation for offline reinforcement learning, which allows for complementing offline data with preference feedback. This framework is crucial for real-world RL applications where direct environment interaction is unsafe or impractical and reward functions are challenging to design manually, yet experts can be queried for their knowledge. (2) In Section 4, we propose a novel offline preference-based RL algorithm that is independent of the specific preference elicitation strategy and recovers a robust policy from an offline dataset and preference feedback. (3) Next, in Section 5, we provide theoretical guarantees on eliciting preferences over samples from the offline dataset, complementing work from Shin et al. [2022]. (4) Then, in Section 6 we propose our own efficient preference elicitation algorithm based on simulated rollouts in a learned environment model. (5) Finally, we establish the theoretical guarantees of our algorithm and demonstrate its empirical efficiency and scalability in different decision-making environments.
@@ 2 Related Work
Our problem setting shares similarities with Offline RL and Preference-based RL, which we summarize below. We position ourselves relative to our closest related works in Table 1.
@@
Framework Offline Efficient Sampling Robustness Guarantees Practical Implementation
@@
OPRL [Shin et al., 2022]  [^1] We demonstrate this in the present work.
@@
Offline RL. Offline Reinforcement Learning has gained significant traction in recent years, as the practicality of training RL agents without environment interaction makes it relevant to real-world applications [Levine et al., 2020]. However, learning from observational data only is a source of bias in the model, as the data may not cover the entire state-action space. Offline RL algorithms therefore output pessimistic policies, which has been shown to minimize suboptimality Jin et al. [2021]. Model-based approaches show particular promise for their sample efficiency [Yu et al., 2020, Kidambi et al., 2020, Rigter et al., 2022, Zhai et al., 2024, Uehara and Sun, 2021]. In this work, we study the setting where reward signals are unavailable and must be estimated by actively querying preference feedback.
Preference-based RL. Rather than accessing numerical reward values for each state-action pair as in traditional online RL, preference-based RL learns the reward model through collecting pairwise preferences over trajectories [Wirth et al., 2017]. Different preference elicitation strategies have been proposed for this framework, generally based on knowing the transition model exactly or on having access to the environment for rollouts [Christiano et al., 2017, Saha et al., 2023, Chen et al., 2022, Lindner et al., 2021, Zhan et al., 2023b, Sadigh et al., 2018, Brown et al., 2020].
Offline Preference-based RL. The development of preference-based RL algorithms based on offline data only is critical to settings where environment interaction is not feasible for safety and efficiency reasons. Still, this framework remains largely unexplored in the literature. While Zhu et al. [2023], Zhan et al. [2023a] demonstrate the value of pessimism in offline preference-based reinforcement learning, they do not consider how to query feedback actively. On the other hand, Shin et al. [2022] propose an empirical comparison of different preference sampling trajectories from an offline trajectories buffer. In Section 5, we provide a theoretical analysis of their approach, then propose an alternative sampling strategy based on simulated trajectory rollouts in Section 6, which enjoys both theoretical and empirical motivation.
@@ 3.1 Preliminaries
Markov Decision Process. We consider the episodic Markov Decision Process (MDP), defined by the tuple ${\mathcal{M}=(\mathcal{S},\mathcal{A},H,T,{R})}$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the episode length, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition function, ${R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. We assume an initial state $s_{0}$, but our analysis could be easily generalized to a fixed initial state distribution. At time $t$, the environment is at state $s_{t}\in\mathcal{S}$ and an agent selects an action $a_{t}\in\mathcal{A}$. The agent then receives a reward ${R}(s_{t},a_{t})$ and the environment transitions to state $s_{t+1}\sim T({}\cdot{}|s_{t},a_{t})$. We describe an agent’s behavior through a policy function $\pi:\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$, such that $\pi(a|s)$ is the probability of taking action $a$ in state $s$. Let $\tau=(s_{0},a_{0},\ldots s_{H},a_{H})$ denote the trajectory of state-action pairs of an interaction episode with the environment. With an abuse of notation, we also write $R(\tau)=\sum_{t}R(s_{t},a_{t})$. Let $d^{\pi}_{T}$ denote the distribution of trajectories induced by rolling out policy $\pi$ in transition model $T$. We denote the expected return of policy $\pi$ as $V^{\pi}_{T,R}=\mathbb{E}_{\tau\sim d^{\pi}_{T}}[R(\tau)]$, and $\pi^{*}=\text{argmax}_{\pi}V^{\pi}_{T,R}$ denotes the optimal policy in $\mathcal{M}$.
Preference-based Reinforcement Learning. Rather than observing numerical rewards at each state and action, we receive preference feedback over trajectories. For a pair of trajectories $(\tau_{1},\tau_{2})$, we obtain binary feedback $o\in\{0,1\}$ about whether $\tau_{1}$ is preferred to $\tau_{2}$. We assume that preference labels follow the Bradley-Terry model [Bradley and Terry, 1952]:
@@
where $\succ$ denotes a preference relationship and $\sigma$ is the sigmoid function. Within this framework, preference elicitation refers to the process of sampling preferences to obtain information about both the preference function and the system dynamics [Wirth et al., 2017].
@@ 3.2 Offline Preference Elicitation
We assume access to an observational dataset of trajectories $\mathcal{D}_{offline}=\{\tau:\tau\sim d^{\pi_{\beta}}_{T}\}$, where $\pi_{\beta}$ is an unknown behavioural policy in $\mathcal{M}$. As in Offline RL, we do not have access to the decision-making environment to observe transition dynamics or rewards under alternative action choices. We assume not to have access to the reward function, but we can query preference feedback from a human to generate a dataset of preferences $\mathcal{D}_{pref}=\{(\tau_{1},\tau_{2},o)\}$.
Optimality Criterion. Based only on our offline dataset $\mathcal{D}_{offline}$, our goal is to recover a policy $\hat{\pi}^{*}$ that minimizes suboptimality in the true environment with as few human preference queries as possible. Let $\pi^{*}_{offline}$ denote the optimal offline policy estimated based on the offline data, with access to the true reward function $R$, and let $\epsilon_{T}$ denote its suboptimality. Since preference elicitation only allows us to estimate the reward function, we do not aim to achieve a suboptimality less than $\epsilon_{T}$. [^2] However, $\epsilon_{T}$ is not formally a lower bound for our problem, as shown in Section A.3. Our objective is then formalized as follows.
@@
Let $\pi^{*}$ be the optimal policy in $\mathcal{M}$ and $\hat{\pi}^{*}$ be the estimated optimal policy based on an offline dataset $\mathcal{D}_{offline}$ and $N_{p}>0$ preference queries. Let $\epsilon_{T}$ be the inherent suboptimality assuming access to the true reward function. We say that a sampling strategy is $(\epsilon,\delta,N_{p})$-correct if for every $\epsilon\geq\epsilon_{T}$, with probability at least $(1-\delta)$, it holds that $V^{\pi^{*}}_{T,R}-V^{\hat{\pi}^{*}}_{T,R}\leq\epsilon$.
@@
The reward function is bounded: $0\leq\tilde{R}(\tau)\leq R_{max}$ for all $\tilde{R}\in\mathcal{F}_{R}$ and all trajectories $\tau$.
@@
Let $\mathcal{F}$ be a class of real functions $f:\mathcal{X}\rightarrow\mathbb{R}$. We say $(l,u)$ is an $\epsilon$-bracket if $l(x)\leq u(x)$ and $\|u(x)-l(x)\|_{1}\leq\epsilon$ for all $x\in\mathcal{X}$. The $\epsilon$-bracketing number of $\mathcal{F}$, denoted $\mathcal{N}_{\mathcal{F}}(\epsilon)$, is the minimal number of $\epsilon$-brackets $(l^{n},u^{n})_{n=1}^{N}$ needed so that for any $f\in\mathcal{F}$, there is a bracket $i\in[N]$ containing it, meaning $l^{i}(x)\leq f(x)\leq u^{i}(x)$ for all $x\in\mathcal{X}$.
Let $\mathcal{N}_{\mathcal{F}_{R}}(\epsilon)$ and $\mathcal{N}_{\mathcal{F}_{T}}(\epsilon)$ denote the $\epsilon$-bracketing numbers of $\mathcal{F}_{R}$ and $\mathcal{F}_{T}$ respectively. This measures the complexity of the function classes [Geer, 2000]. For instance, with linear rewards $\mathcal{F}_{R}\vcentcolon=\{R:\tau\rightarrow\theta^{T}\phi(\tau)\}$, the $\epsilon$-bracket number is bounded by $\mathcal{N}_{\mathcal{F}_{R}}(\epsilon)\leq\mathcal{O}(d\log\frac{BR}{\epsilon})$, where $\|\theta\|_{2}\leq B$ and $\|\phi(\tau)\|_{2}\leq R\>\forall\tau\in\mathcal{T}$, and $d$ is the dimension of the feature space [Zhan et al., 2023a].
@@
$C_{T}(\mathcal{F}_{T},\pi^{*})=\sup_{\tilde{T}\in\mathcal{F}_{T}}\left[\frac{\mathbb{E}_{(s,a)\sim d^{\pi^{*}}_{T}}[|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)|]}{\sqrt{\mathbb{E}_{(s,a)\sim\mathcal{D}_{offline}}[|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)|^{2}]}}\right]$
The concentrability coefficient measures the coverage of the optimal policy in the offline dataset. Note that $C_{T}$ is upper-bounded by the density-ratio coefficient: $C_{T}(\mathcal{F}_{T},\pi^{*})\leq\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}{d^{\pi^{*}}_{T}(s,a)}/{d^{\pi_{\beta}}_{T}(s,a)}$, where $\pi_{\beta}$ is the behavioural policy underlying $\mathcal{D}_{offline}$.
@@ 4 Offline Preference-based RL with Preference Elicitation
In this section, we propose a general framework for offline preference-based reinforcement learning. The next two sections propose two different preference elicitation strategies. As learning must be carried out in two stages, with environment dynamics based on $\mathcal{D}_{offline}$ and reward learning on $\mathcal{D}_{pref}$, we adopt a model-based approach which we summarize in Algorithm 1.
@@
$\hat{T}\quad$ $=\text{argmax}_{\tilde{T}\in\mathcal{F}_{T}}\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{offline}}\left[\log\tilde{T}(s^{\prime}|s,a)\right],$
$u_{T}(s,a)$ $=\max_{\tilde{T}_{1},\tilde{T}_{2}\in\mathcal{T}}|\tilde{T}_{1}(\cdot|s,a)-\tilde{T}_{2}(\cdot|s,a)|\cdot{R_{max}},$
where $\mathcal{T}=\{\tilde{T}\in\mathcal{F}_{T}\>|\>\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{offline}}\left[\log{\hat{T}(s^{\prime}|s,a)}/{\tilde{T}(s^{\prime}|s,a)}\right]\leq\beta_{T}\}$, defining a confidence set over the MLE estimate, and $\beta_{T}$ is a hyperparameter. In a practical implementation, this can be achieved by training an ensemble of models on different data bootstraps [Lakshminarayanan et al., 2017].
@@
1:Observational trajectories dataset $\mathcal{D}_{offline}$. Significance $\delta\in(0,1)$, preference budget $N_{p}$.
@@
3:Estimate $\hat{T}$ and $u_{T}$ via maximum likelihood over the observational data $\mathcal{D}_{offline}$.
4:$\mathcal{D}_{pref}\leftarrow\emptyset$.
@@
8:    $\mathcal{D}_{pref}\leftarrow\mathcal{D}_{pref}\cup\{(\tau_{1},\tau_{2},o)\}$.
9:    Estimate $\hat{R}$ and $u_{R}$ via maximum likelihood over the preference data $\mathcal{D}_{pref}$.
@@
Iterative Preference Elicitation and Reward Learning. As with the transition model, our algorithm estimates the reward function $\hat{R}$ and its uncertainty function through maximum likelihood over iteratively collected preference data $\mathcal{D}_{pref}$:
$\hat{R}\quad$ $=\text{argmax}_{\tilde{R}\in\mathcal{F}_{R}}\mathbb{E}_{(\tau_{1},\tau_{2},o)\sim\mathcal{D}_{pref}}\left[o\log P_{\tilde{R}}(\tau_{1}\succ\tau_{2})+(1-o)\log P_{\tilde{R}}(\tau_{2}\succ\tau_{1})\right],$
$u_{R}(\tau)$ $=\max_{\tilde{R}_{1},\tilde{R}_{2}\in\mathcal{R}}|\tilde{R}_{1}(\tau)-\tilde{R}_{2}(\tau)|,$
where =$\mathcal{R}=\{\tilde{R}\in\mathcal{F}_{R}\>|\>\mathbb{E}_{(\tau_{1},\tau_{2},o)\sim\mathcal{D}_{pref}}\left[\log{P_{\hat{R}}(\tau_{1}\succ\tau_{2})}/{P_{\tilde{R}}(\tau_{1}\succ\tau_{2})}\right]\leq\beta_{R}\}$ defines the confidence set and $\beta_{R}$ is a hyperparameter. We also define preference uncertainty between two trajectories $\tau_{1},\tau_{2}$ as $u_{P_{R}}(\tau_{1},\tau_{2})=\max_{\tilde{R}_{1},\tilde{R}_{2}\in\mathcal{R}}|P_{\tilde{R}_{1}}(\tau_{1}\succ\tau_{2})-P_{\tilde{R}_{2}}(\tau_{1}\succ\tau_{2})|$.
The choice of trajectory sampling strategy for preference elicitation in line 4 is critical to efficiently obtaining an $\epsilon$-optimal policy. We present two possible strategies in Sections 5 and 6. Note that by focusing on sample efficiency as in prior work on preference elicitation [Chen et al., 2022], we do not necessarily optimize for computational efficiency; this could be improved by collecting preferences in batches to reduce the number of reward training loops.
@@
$\hat{\pi}^{*}=\text{argmax}_{\pi\in\Pi}\min_{\tilde{T}\in\mathcal{T},\tilde{R}\in\mathcal{R}}V^{\pi}_{\tilde{T},\tilde{R}}.$
This analysis provides a worst-case robustness guarantee when considering well-calibrated confidence intervals, as detailed in Sections 5.1 and 6.1. For a practical implementation of our algorithm, we penalize the reward function by the uncertainty as in model-based offline RL methods [Yu et al., 2020, Chang et al., 2021]. Our optimal robust policy therefore maximizes the following objective:
$\hat{\pi}^{*}=\text{argmax}_{\pi\in\Pi}\mathbb{E}_{\tau\sim d^{\pi}_{\hat{T}}}[\hat{R}(\tau)-u_{R}(\tau)-u_{T}(\tau)].$ (2)
We show in Section A.2 that this is indeed a lower bound of the true value function. This objective allows for controlling the degree of conservatism in practice through the width of the confidence intervals used to determine $u_{R}$ and $u_{T}$.

### v2

@@ Abstract
Applying reinforcement learning (RL) to real-world problems is often made challenging by the inability to interact with the environment and the difficulty of designing reward functions. Offline RL addresses the first challenge by considering access to an offline dataset of environment interactions labeled by the reward function. In contrast, Preference-based RL does not assume access to the reward function and learns it from preferences, but typically requires an online interaction with the environment. We bridge the gap between these frameworks by exploring efficient methods for acquiring preference feedback in a fully offline setup. We propose Sim-OPRL, an offline preference-based reinforcement learning algorithm, which leverages a learned environment model to elicit preference feedback on simulated rollouts. Drawing on insights from both the offline RL and the preference-based RL literature, our algorithm employs a pessimistic approach for out-of-distribution data, and an optimistic approach for acquiring informative preferences about the optimal policy. We provide theoretical guarantees regarding the sample complexity of our approach, dependent on how well the offline data covers the optimal policy. Finally, we demonstrate the empirical performance of Sim-OPRL in various environments.
@@ 1 Introduction
Our contributions are the following: (1) In Section 3, we first formalize the new problem setting of preference elicitation for offline reinforcement learning, which allows for complementing offline data with preference feedback. This framework is crucial for real-world RL applications where direct environment interaction is unsafe or impractical and reward functions are challenging to design manually, yet experts can be queried for their knowledge. (2) In Section 5, we provide theoretical guarantees on eliciting preference feedback over samples from an offline dataset, complementing work of Shin et al. [2022]. (3) Next, in Section 6, we propose our own efficient preference elicitation algorithm based on simulated rollouts in a learned environment model, and establish its improved theoretical guarantees. (4) Finally, we develop a practical implementation of our algorithm and demonstrate its empirical efficiency and scalability across various decision-making environments.
@@ 2 Related Work
Our problem setting shares similarities with Offline RL and Preference-based RL, which we summarize below. We position ourselves relative to our closest related works in Table 1 and extend our discussion in Appendix B.
@@
Framework Offline Policy-Based Sampling Robustness Guarantees Practical Implementation
@@
OPRL [Shin et al., 2022]
@@
Offline RL. Offline Reinforcement Learning has gained significant traction in recent years, as the practicality of training RL agents without environment interaction makes it relevant to real-world applications [Levine et al., 2020]. However, learning from observational data only is a source of bias in the model, as the data may not cover the entire state-action space. Offline RL algorithms therefore output pessimistic policies, which has been shown to minimize suboptimality [Jin et al., 2021]. Model-based approaches show particular promise for their sample efficiency [Yu et al., 2020, Kidambi et al., 2020, Rigter et al., 2022, Zhai et al., 2024, Uehara and Sun, 2021]. In this work, we study the setting where reward signals are unavailable and must be estimated by actively querying preference feedback.
Preference-based RL. Rather than accessing numerical reward values for each state-action pair as in traditional online RL, preference-based RL learns the reward model through collecting pairwise preferences over trajectories [Wirth et al., 2017]. Different preference elicitation strategies have been proposed for this framework, generally based on knowing the transition model exactly or on having access to the environment for rollouts [Christiano et al., 2017, Saha et al., 2023, Chen et al., 2022, Lindner et al., 2021, Zhan et al., 2023b, Sadigh et al., 2018, Brown et al., 2020]. Prior theoretical and empirical work [Lindner et al., 2021, Chen et al., 2022] show that, in this setting, the most efficient preference elicitation strategy is to actively reduce the set of candidate optimal policies, rather maximize information gain on the reward function – our theoretical and empirical results reach the same conclusion for the offline setting.
Offline Preference-based RL. The development of preference-based RL algorithms based on offline data only is critical to settings where environment interaction is not feasible for safety and efficiency reasons. Still, this framework remains largely unexplored in the literature. While Zhu et al. [2023], Zhan et al. [2023a] demonstrate the value of pessimism in offline preference-based reinforcement learning, they do not consider how to query feedback actively. On the other hand, Shin et al. [2022] propose an empirical comparison of different preference sampling trajectories from an offline trajectories buffer. In Section 5, we provide a theoretical analysis of their approach, then propose an alternative sampling strategy based on simulated trajectory rollouts in Section 6, which benefits from both theoretical motivation and superior empirical performance.
@@ 3.1 Preliminaries
Markov Decision Process. We consider the episodic Markov Decision Process (MDP), defined by the tuple ${\mathcal{M}=(\mathcal{S},\mathcal{A},H,T,{R})}$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $H$ is the episode length, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta_{\mathcal{S}}$ is the transition function, ${R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function. We assume an initial state $s_{0}$, but our analysis generalizes to a fixed initial state distribution. At time $t$, the environment is at state $s_{t}\in\mathcal{S}$ and an agent selects an action $a_{t}\in\mathcal{A}$. The agent then receives a reward ${R}(s_{t},a_{t})$ and the environment transitions to state $s_{t+1}\sim T({}\cdot{}|s_{t},a_{t})$. We describe an agent’s behavior through a policy function $\pi:\mathcal{S}\rightarrow\Delta_{\mathcal{A}}$, such that $\pi(a|s)$ is the probability of taking action $a$ in state $s$. Let $\tau=(s_{0},a_{0},\ldots s_{H},a_{H})$ denote the trajectory of state-action pairs of an interaction episode with the environment. With an abuse of notation, we also write $R(\tau)=\sum_{t}R(s_{t},a_{t})$. Let $d^{\pi}_{T}$ denote the distribution of trajectories (or state-action pairs, overloading notation) induced by rolling out policy $\pi$ in transition model $T$. We denote the expected return of policy $\pi$ as $V^{\pi}_{T,R}=\mathbb{E}_{\tau\sim d^{\pi}_{T}}[R(\tau)]$, and $\pi^{*}=\text{argmax}_{\pi}V^{\pi}_{T,R}$ denotes the optimal policy in $\mathcal{M}$.
Preference-based Reinforcement Learning. Rather than observing rewards for each state and action, we receive preference feedback over trajectories. For a pair of trajectories $(\tau_{1},\tau_{2})$, we obtain binary feedback $o\in\{0,1\}$ about whether $\tau_{1}$ is preferred to $\tau_{2}$. We define the preference function $P_{R}$ and assume that preference labels follow the Bradley-Terry model [Bradley and Terry, 1952]:
@@
where $\succ$ denotes a preference relationship and $\sigma$ is a monotonous increasing link function. Within this framework, preference elicitation refers to the process of sampling preferences to obtain information about both the preference function and the system dynamics [Wirth et al., 2017].
@@ 3.2 Offline Preference Elicitation
We assume access to an observational dataset of trajectories $\mathcal{D}_{\textrm{offline}}=\{\tau:\tau\sim d^{\pi_{\beta}}_{T}\}$, where $\pi_{\beta}$ is an unknown behavioural policy in $\mathcal{M}$. As in Offline RL, we do not have access to the environment to observe transition dynamics or rewards under alternative action choices. We assume not to have access to the reward function, but we can query preference feedback from a human to generate a dataset of preferences $\mathcal{D}_{\textrm{pref}}=\{(\tau_{1},\tau_{2},o)\}$.
Optimality Criterion. Based only on our offline dataset $\mathcal{D}_{\textrm{offline}}$, our goal is to recover a policy $\hat{\pi}^{*}$ that minimizes suboptimality in the true environment with as few human preference queries as possible. Let $\pi^{*}_{\textrm{offline}}$ denote the optimal offline policy estimated based on the offline data, with access to the true reward function $R$, and let $\epsilon_{T}$ denote its suboptimality. Since preference elicitation only allows us to estimate the reward function, we do not aim to achieve a suboptimality less than $\epsilon_{T}$ – although this is not formally a lower bound for our problem, as shown in Section A.3. Our objective is then formalized as follows.
@@
Let $\pi^{*}$ be the optimal policy in $\mathcal{M}$ and $\hat{\pi}^{*}$ be the estimated optimal policy based on an offline dataset $\mathcal{D}_{\textrm{offline}}$ and $N_{p}>0$ preference queries. Let $\epsilon_{T}$ be the inherent suboptimality assuming access to the true reward function. We say that a sampling strategy is $(\epsilon,\delta,N_{p})$-correct if for every $\epsilon\geq\epsilon_{T}$, with probability at least $(1-\delta)$, it holds that $V^{\pi^{*}}_{T,R}-V^{\hat{\pi}^{*}}_{T,R}\leq\epsilon$.
@@
The reward function is bounded: $0\leq\tilde{R}(\tau)\leq R_{\textup{max}}$ for all $\tilde{R}\in\mathcal{F}_{R}$ and all trajectories $\tau$.
@@
Let $\mathcal{F}$ be a class of real functions $f:\mathcal{X}\rightarrow\mathbb{R}$. We say $(l,u)$ is an $\epsilon$-bracket if $l(x)\leq u(x)$ and $|u(x)-l(x)|_{1}\leq\epsilon$ for all $x\in\mathcal{X}$. The $\epsilon$-bracketing number of $\mathcal{F}$, denoted $\mathcal{N}_{\mathcal{F}}(\epsilon)$, is the minimal number of $\epsilon$-brackets $(l^{n},u^{n})_{n=1}^{N}$ needed so that for any $f\in\mathcal{F}$, there is a bracket $i\in[N]$ containing it, meaning $l^{i}(x)\leq f(x)\leq u^{i}(x)$ for all $x\in\mathcal{X}$.
Let $\mathcal{N}_{\mathcal{F}_{R}}(\epsilon)$ and $\mathcal{N}_{\mathcal{F}_{T}}(\epsilon)$ denote the $\epsilon$-bracketing numbers of $\mathcal{F}_{R}$ and $\mathcal{F}_{T}$ respectively. This measures the complexity of the function classes [Geer, 2000]. For instance, with linear rewards $\mathcal{F}_{R}\vcentcolon=\{R:\tau\rightarrow\theta^{T}\phi(\tau)\}$, the $\epsilon$-bracket number is bounded by $\log\mathcal{N}_{\mathcal{F}_{R}}(\epsilon)\leq\mathcal{O}(d\log\frac{BG}{\epsilon})$, where $\|\theta\|_{2}\leq B$ and $\|\phi(\tau)\|_{2}\leq G\>\forall\tau\in\mathcal{T}$, and $d$ is the dimension of the feature space [Zhan et al., 2023a].
@@
$C_{T}(\mathcal{F}_{T},\pi^{*})=\sup_{\tilde{T}\in\mathcal{F}_{T}}\left[\frac{\mathbb{E}_{(s,a)\sim d^{\pi^{*}}_{T}}[|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)|_{1}]}{\sqrt{\mathbb{E}_{(s,a)\sim\mathcal{D}_{\textup{offline}}}[|T(\cdot|s,a)-\tilde{T}(\cdot|s,a)|_{1}^{2}]}}\right],$
The concentrability coefficient measures the coverage of the optimal policy in the offline dataset. Note that $C_{T}$ is upper-bounded by the density-ratio coefficient: $C_{T}(\mathcal{F}_{T},\pi^{*})\leq\sup_{(s,a)\in\mathcal{S}\times\mathcal{A}}{d^{\pi^{*}}_{T}(s,a)}/{d^{\pi_{\beta}}_{T}(s,a)}$, where $\pi_{\beta}$ is the behavioural policy underlying $\mathcal{D}_{\textrm{offline}}$.
@@ 4 Offline Preference-based RL with Preference Elicitation
In this section, we propose a general framework for offline preference-based reinforcement learning. The next two sections propose two different preference elicitation strategies. As learning must be carried out in two distinct stages, with environment dynamics based on $\mathcal{D}_{\textrm{offline}}$ and reward learning on $\mathcal{D}_{\textrm{pref}}$, we adopt a model-based approach which we summarize in Algorithm 1.
@@
$\hat{T}\quad$ $=\text{argmax}_{\tilde{T}\in\mathcal{F}_{T}}\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\textrm{offline}}}\left[\log\tilde{T}(s^{\prime}|s,a)\right],$
$u_{T}(s,a)$ $=\max_{\tilde{T}_{1},\tilde{T}_{2}\in\mathcal{T}}|\tilde{T}_{1}(\cdot|s,a)-\tilde{T}_{2}(\cdot|s,a)|_{1}\cdot{R_{\textrm{max}}},$
where $\mathcal{T}=\{\tilde{T}\in\mathcal{F}_{T}\>|\>\mathbb{E}_{(s,a,s^{\prime})\sim\mathcal{D}_{\textrm{offline}}}\left[\log\left({\hat{T}(s^{\prime}|s,a)}/{\tilde{T}(s^{\prime}|s,a)}\right)\right]\leq\beta_{T}\}$, defining a confidence set over the maximum likelihood estimator (MLE), and $\beta_{T}$ is a margin hyperparameter.
@@
1:Observational trajectories dataset $\mathcal{D}_{\textrm{offline}}$. Significance $\delta\in(0,1)$, preference budget $N_{p}$.
@@
3:Estimate $\hat{T}$ and $u_{T}$ via maximum likelihood over the observational data $\mathcal{D}_{\textrm{offline}}$.
4:$\mathcal{D}_{\textrm{pref}}\leftarrow\emptyset$.
@@
8:     $\mathcal{D}_{\textrm{pref}}\leftarrow\mathcal{D}_{\textrm{pref}}\cup\{(\tau_{1},\tau_{2},o)\}$.
9:     Estimate $\hat{R}$ and $u_{R}$ via maximum likelihood over the preference data $\mathcal{D}_{\textrm{pref}}$.
@@
Iterative Preference Elicitation and Reward Learning. As with the transition model, our algorithm estimates the reward function $\hat{R}$ and its uncertainty function through maximum likelihood over iteratively collected preference data $\mathcal{D}_{\textrm{pref}}$:
$\hat{R}\quad$ $=\text{argmax}_{\tilde{R}\in\mathcal{F}_{R}}\mathbb{E}_{(\tau_{1},\tau_{2},o)\sim\mathcal{D}_{\textrm{pref}}}\left[o\log P_{\tilde{R}}(\tau_{1}\succ\tau_{2})+(1-o)\log P_{\tilde{R}}(\tau_{2}\succ\tau_{1})\right],$
$u_{R}(\tau)$ $=\max_{\tilde{R}_{1},\tilde{R}_{2}\in\mathcal{R}}|\tilde{R}_{1}(\tau)-\tilde{R}_{2}(\tau)|_{1},$
where =$\mathcal{R}=\{\tilde{R}\in\mathcal{F}_{R}\>|\>\mathbb{E}_{(\tau_{1},\tau_{2},o)\sim\mathcal{D}_{\textrm{pref}}}\left[\log\left({P_{\hat{R}}(\tau_{1}\succ\tau_{2})}/{P_{\tilde{R}}(\tau_{1}\succ\tau_{2})}\right)\right]\leq\beta_{R}\}$ defines the confidence set and $\beta_{R}$ is a hyperparameter. We also define preference uncertainty between two trajectories $\tau_{1},\tau_{2}$:
$u_{P_{R}}(\tau_{1},\tau_{2})=\max_{\tilde{R}_{1},\tilde{R}_{2}\in\mathcal{R}}|P_{\tilde{R}_{1}}(\tau_{1}\succ\tau_{2})-P_{\tilde{R}_{2}}(\tau_{1}\succ\tau_{2})|_{1}.$ (2)
The choice of trajectory sampling strategy for preference elicitation in line 4 (Algorithm 1) is critical to efficiently obtaining an $\epsilon$-optimal policy. We present two possible strategies in Sections 5 and 6.
@@
$\hat{\pi}^{*}=\text{argmax}_{\pi\in\Pi}\min_{\tilde{T}\in\mathcal{T},\tilde{R}\in\mathcal{R}}V^{\pi}_{\tilde{T},\tilde{R}}.$ (3)
Based on this objective, we define the pessimistic transition and reward models as follows: $\hat{T}_{\text{inf}},\hat{R}_{\text{inf}}=\text{argmin}_{\tilde{T}\in\mathcal{T},\tilde{R}\in\mathcal{R}}\max_{\pi\in\Pi}V^{\pi}_{\tilde{T},\tilde{R}}$. Our analysis provides a worst-case robustness guarantee when considering well-calibrated confidence intervals, as detailed in Sections 5.1 and 6.1. In other words, following prior work [Chen et al., 2022, Zhan et al., 2023a], our theoretical analysis assumes that modeling elements can be identified with no optimization error. We then complement this algorithmic framework with a flexible practical implementation in Section 6.3.